# Reservoir Computing as a Language Model

### v2

@@ VII Results
Figure 3 presents training (top row) and testing (bottom row) losses for the five model sizes listed in Table 1, ordered from smallest (turquoise) to largest (grey). Model sizes range from approximately 15,000 to 155,000 trainable parameters.
@@
Analyzing the loss curves for the Transformer, the classical Reservoir Computer (RC), and the Attention-Enhanced Reservoir Computer (AERC), we observe that loss variance across different model sizes is more pronounced in the RC and AERC than in the Transformer. As expected, the Transformer achieves the best overall performance, with a minimum test loss of 1.67 for the largest configuration. However, we note that the classical RC achieves a test loss as low as 2.01, and the AERC reaches an intermediate value of 1.73—indicating that both RC-based models are competitive despite their relative simplicity.
@@
An important direction for further research is to investigate scaling laws for reservoir computing in the context of large language model (LLM) prediction. These results would be consistent with known scaling laws of transformer architectures, where performance improves predictability with increased model size and dataset scale [26]. Understanding how performance scales with reservoir size and model capacity could offer valuable insights into the potential of RC-based architectures in more complex sequence modeling tasks. As computational resources improve, we plan to explore the behavior of both classical and attention-enhanced reservoir computers at larger scales, assessing their viability as alternatives to conventional deep learning models.
VII.1 Evaluation Metrics and Analysis
@@ VII.1.1 N-gram Overlap
VII.1.2 Mathematical Formulation
Given a generated text $G=\{g_{1},g_{2},\dots,g_{M}\}$ and the reference text $R=\{r_{1},r_{2},\dots,r_{N}\}$ we extracted the n-grams as:
$G_{n}=\left\{(g_{i},g_{i+1},\dots,g_{i+n-1})\mid 1\leq i\leq M-n+1\right\},$
$R_{n}=\left\{(r_{j},r_{j+1},\dots,r_{j+n-1})\mid 1\leq j\leq N-n+1\right\}.$
The overlap-n metric is then given by:
[…]
This metric assesses how well the model captures the n-gram patterns present in the reference data. A higher overlap indicates that the generated text shares more common sequences with the reference, suggesting better generalization to the learned patterns.
VII.2 N-Gram Similarity
The results of the 7-gram and 8-gram performance are shown in Fig. 4. The evaluation metrics reveal that the classic reservoir approach tends to have the lowest overlap and thus the lowest generalization. Transformers and Attention-Enhanced Reservoirs demonstrate on par diversity and distributional alignment, making them well-suited for tasks like text generation. Their ability to maintain high performance underscores their superior performance, due to the inclusion of neural networks.
@@ VII.2 N-Gram Similarity
The attention-enhanced reservoir improves upon the traditional model by incorporating dynamic readout weights, thereby enhancing its ability to capture more complex patterns without incurring the full computational burden of transformers. This architecture serves as a middle ground, offering improved performance over the traditional reservoir, while offering reducing complexity in the neural network architecture.
VII.3 Inference Time Comparison
Figure 5: Training and inference time for a longer text generation over the number of trainable parameters for the Transformer, Reservoir, and Attention-Enhanced Reservoir Computer. The numbers above the line are from a fit applied via $y=\alpha\log_{10}(x)$, where $x$ is the number of trainable parameters, $y$ is the training time and $\alpha$ the slope.
@@ VII.3 Inference Time Comparison
In Figure 5, solid lines indicate training time, while dashed lines represent inference time. The color scheme distinguishes between the three model types: classic reservoir (blue), attention-enhanced reservoir (orange), and transformer (green, red, and purple). It is evident that reservoir-based models exhibit substantially lower computational times, and their scalability with respect to parameter count is more favorable than that of transformers. The numbers above the line are from a fit applied via $y=\alpha\log_{10}(x)$, where $x$ is the number of trainable parameters, $y$ is the training time and $\alpha$ the slope. From the graphs itself and the slope $\alpha$ we can see that reservoir based training and inference time is by magnitudes of order faster. This performance advantage arises because the reservoir handles the computationally expensive sequence processing, whereas in transformers, attention mechanisms scale quadratically with sequence length. These results suggest that using a reservoir for data preprocessing can significantly accelerate language model performance without incurring the high time cost associated with full transformer architectures.
VIII Conclusion
@@
We presented a unified framework for comparing classical reservoir computing (with a static linear readout), an attention-enhanced reservoir (with a dynamic readout), and transformer architectures on a character-level Shakespeare corpus. By systematically varying model sizes and measuring cross-entropy, N-gram Overlap, and time costs, we revealed how each architecture scales in performance and resource usage. Transformers excel overall, solidifying their state-of-the-art machine learning status, though they are more demanding in terms of computation time due to the quadratic complexity of the sequence length. Traditional reservoirs, while limited in their ability to capture complex dependencies, show surprising good results for their simple buildup. Attention-enhanced reservoirs strike a balance between efficiency and performance, outperforming traditional reservoirs and even getting close to the performance of Transformer models while having a negotiable complexity depending on the sequence length. The code used for this study is available at https://github.com/fekoester/Shakespeare_Res.
@@ Acknowledgements.
This study was supported in part by JSPS KAKENHI (JP22H05195, JP25H01129) and JST CREST (JPMJCR24R2) in Japan.

### v3

@@ III.1 Comparison of Training and Testing Loss
To quantitatively assess model performance, we report training and testing losses across epochs and data shards for all five model sizes described in Table 1. The results are visualized in Figure 2. Figure 2 top row presents training and bottom row testing losses for the five model sizes listed in Table 1, ordered from smallest (turquoise) to largest (grey), where the first column shows the transformer, the second column the reservoir, and the third column the AERC. Model sizes range from approximately 15,000 to 155,000 trainable parameters.
@@
Analyzing the loss curves for the Transformer, the classical Reservoir Computer (RC), and the Attention-Enhanced Reservoir Computer (AERC), we observe that loss variance across different model sizes is more pronounced in the RC and AERC than in the Transformer. As expected, the Transformer achieves the best overall performance, with a minimum test loss of 1.67 for the largest configuration. However, we note that the classical RC achieves a test loss as low as 1.81, and the AERC reaches an intermediate value of 1.73—indicating that both RC-based models are competitive despite their relative simplicity.
@@
Figure 4: Training (solid) and inference (dashed) time for a longer text generation over the number of trainable parameters for the Transformer, Reservoir, and Attention-Enhanced Reservoir Computer. The numbers above the line are from a fit applied via $y=\alpha\log_{10}(x)$, where $x$ is the number of trainable parameters, $y$ is the training time and $\alpha$ the slope.
III.2 Evaluation Metrics and Analysis
@@
InputtokensTokenembeddingPositionalencodingEmbedding ++positionLAERC block 1LAERC block 2LAERC block $L$ $L$ stacked blocksFinalLayerNormLinearprojectionSoftmax(next token)
(a) LAERC language model.
@@
(b) Single LAERC block.
Figure 5: Overview of the layered attention-enhanced reservoir computing (LAERC) architecture: (a) the full language model with $L$ stacked reservoir blocks and (b) the internal structure of a single block combining a fixed reservoir, gating, and ReZero feed-forward refinement. The dashed lines indicate residual connections, enabling the LAERC to backpropagate past the fixed unknown reservoir.
$\text{Overlap-n}=\frac{\left|\left\{G_{n}\cap R_{n}\right\}\right|}{\left|G_{n}\right|},$ (8)
@@
This metric assesses how well the model captures the n-gram patterns present in the reference data. A higher overlap indicates that the generated text shares more common sequences with the reference.
III.3 N-Gram Similarity
The results of the 7-gram and 8-gram performance are shown in Fig. 3. The evaluation metrics reveal that the classic reservoir approach tends to have the lowest overlap and thus the lowest generalization. Transformers and Attention-Enhanced Reservoirs demonstrate on par diversity and distributional alignment. Their ability to maintain higher performance underscores their superior performance, due to the inclusion of neural networks.
@@ III.3 N-Gram Similarity
The attention-enhanced reservoir improves upon the traditional model by incorporating dynamic readout weights, thereby enhancing its ability to capture more complex patterns.
Figure 6: Smoothed training loss as a function of the global optimization step for transformer baselines and layered reservoir (ReservoirFFN) language models trained on OpenWebText. Each curve corresponds to a different embedding dimension and thus a different total parameter count.
III.4 Inference Time Comparison
@@
Furthermore, although transformer architectures benefit from high degrees of parallelization in digital hardware, the sequential nature of reservoir dynamics does not impose a practical bottleneck when implemented on physical substrates. Experimental demonstrations using photonic and analog reservoirs show operational speeds over GHz the internal state evolution occurs effectively instantaneously compared to the digitally executed postprocessing stages [3, 28]. Therefore, physical reservoirs can be regarded as practically parallel from the perspective of overall system throughput. In contrast, transformer attention layers remain constrained by digital memory bandwidth and quadratic scaling with respect to sequence length.
@@
In Figure 4, solid lines indicate training time, while dashed lines represent inference time. The color scheme distinguishes between the three model types: classic reservoir (orange), attention-enhanced reservoir (blue), and transformer (green, red, and purple). It is evident that reservoir-based models exhibit substantially lower computational times, and their scalability with respect to parameter count is more favorable than that of transformers. The numbers above the line are from a fit applied via $y=\alpha\log_{10}(x)$, where $x$ is the number of trainable parameters, $y$ is the training time and $\alpha$ the slope. From the graphs itself and the slope $\alpha$ we can see that reservoir based training and inference time is by magnitudes of order faster. This performance advantage arises because the reservoir handles the computationally expensive sequence processing, whereas in transformers, attention mechanisms scale quadratically with sequence length. These results suggest that using a reservoir for data preprocessing can significantly accelerate language model performance without incurring the high time cost associated with full transformer architectures.
IV Layered Reservoir Language Model
While the previous sections focused on single-reservoir and attention-enhanced reservoir architectures for character-level language modeling with smaller models for edge-computing scenarios, we now extend this framework to a layered reservoir language model designed for large-scale token-level training. The central idea is to replace the quadratic-complexity self-attention layers of transformers with fixed recurrent reservoirs, while retaining lightweight trainable components for representation refinement, gating, and output projection.
We refer to this architecture as Layered Attention-Enhanced Reservoir Computing (LAERC). Each layer contains a fixed, untrained recurrent reservoir that performs temporal mixing, combined with trainable feed-forward and gating components that adaptively integrate reservoir-induced features into the model’s representation. Residual connections allow gradients to bypass the fixed reservoirs, enabling end-to-end training using standard backpropagation despite the reservoirs themselves being untrainable.
IV.1 Architecture
A token sequence is first mapped to a sequence of continuous embeddings in the same manner as standard transformer language models. These embeddings are then processed by a stack of $L$ identical reservoir blocks, each operating sequentially. A self-contained technical definition can be found in Sec. A.
Each LAERC block consists of four conceptual components: a fixed recurrent reservoir, a small trainable MLP that projects reservoir states back into the model dimension, an input-dependent mixing gate, and a feed-forward refinement stage. The overall structure is illustrated in Fig. 5.
Figure 7: Best (minimum) training loss achieved by transformer and layered reservoir (ReservoirFFN) language models as a function of the number of trainable parameters (log scale). Straight lines show least-squares fits of the form $\mathcal{L}_{\min}=a\log_{10}N_{\mathrm{param}}+b$, revealing power-law-like scaling for both architectures, with a steeper slope for transformers.
Fixed reservoir.
Within each block, the incoming representation is first normalized and then passed through a single-layer recurrent reservoir with a tanh nonlinearity. The reservoir parameters are randomly initialized once and remain fixed throughout training. Following standard echo-state principles, the recurrent dynamics are scaled to ensure stable temporal behavior, with the effective memory depth increasing mildly with layer index to provide a hierarchy of time scales across depth.
The reservoir state is normalized before further processing. No gradients are propagated through the reservoir itself; instead, learning occurs exclusively in the surrounding trainable components.
Reservoir MLP and scaling.
To map the reservoir state back to the model dimension, each block contains a small trainable MLP. This MLP is intentionally narrow, serving primarily as a readout that extracts useful features from the high-dimensional reservoir dynamics. A learned scalar controls the overall contribution of the reservoir path, allowing the model to adaptively adjust the influence of reservoir-induced features during training. This mechanism stabilizes optimization by preventing overly strong reservoir effects early in training.
Input-dependent mixing gate.
The block combines the original representation and the reservoir-refined representation using an input-dependent gating mechanism. The gate operates at the feature level and determines, for each token, how strongly the model should rely on reservoir-induced updates versus the original signal given by a residual connection. This allows the architecture to interpolate smoothly between purely feed-forward behavior and strong temporal mixing, depending on context.
Feed-forward refinement
Following the mixing step, each block applies a standard position-wise feed-forward network as is typical for a transformer setup. The feed forward dimension is set to 4 times the embedding dimensions.
Output layer.
After passing through all $L$ blocks, the final representation is normalized and projected back into vocabulary space to produce next-token logits. The model is trained using the standard next-token cross-entropy objective over BPE-tokenized text, identical to conventional transformer language models.
Overall, the resulting Reservoir Feed-Forward Network (FFN) replaces transformer self-attention with fixed recurrent reservoirs, while preserving the depth, normalization structure, and feed-forward refinement characteristic of modern LLM architectures. All trainable parameters reside in the embeddings, gating networks, small MLPs, feed-forward layers, and output projection.
IV.2 Training setup on large-scale token data
To evaluate the layered reservoir architecture in a realistic LLM setting, we train it on a byte-pair encoded corpus derived from OpenWebText using a GPT-2-style vocabulary and a fixed context length. The data are stored as memory-mapped token streams for efficient sampling of contiguous training sequences.
We consider a family of models in which the embedding dimension, reservoir size, and number of layers are scaled jointly to produce parameter counts comparable to compact transformer baselines. All models are trained using the same hyperparameters. See Appendix A for details.
IV.3 Results: optimization behavior and scaling laws
Figure 6 shows the training loss as a function of optimization step for both transformer baselines and LAERC models across multiple model sizes. As expected, transformers achieve lower absolute loss at a fixed parameter budget. However, the layered reservoir models exhibit stable and well-behaved optimization dynamics across all evaluated configurations, with performance consistently improving as model capacity increases.
To quantify how the layered reservoir models scale with size and compare with the transformer scaling [11], we extract, for each configuration, the best (minimum) training loss attained during optimization and plot it against the number of trainable parameters on a logarithmic scale, as shown in Fig. 7. Both transformers and LAERC models obey approximate power-law scaling relationships of the form
$\mathcal{L}_{\min}(N_{\mathrm{param}})\approx a\log_{10}N_{\mathrm{param}}+b,$ (9)
where $N_{\mathrm{param}}$ denotes the total number of trainable parameters and $(a,b)$ are fitting coefficients. For the transformer baselines we obtain a slope $a_{\mathrm{TF}}\approx-1.72$, which is a high scaling law value raising from the smaller parameter regime investigated here, while the LAERC models exhibit a weaker but still substantial slope $a_{\mathrm{Res}}\approx-0.81$.
While these results indicate systematic scaling behavior, the present LAERC implementation should be viewed as a baseline rather than an optimized architecture. Several design choices (e.g., reservoir parameterization, coupling topology, and training protocol) were not exhaustively tuned in this study. Future work can therefore focus on architectural and optimization refinements to assess the extent to which the observed scaling trends can be strengthened.
V Conclusion
@@
We presented a unified framework for comparing classical reservoir computing (with a static linear readout), attention-enhanced reservoir architectures (with a dynamic readout), and transformer architectures on a small character-level Shakespeare corpus for edge computing with low computation. By systematically varying model sizes and evaluating cross-entropy, N-gram overlap, and training/inference time cost, we demonstrated how each architecture scales in performance and computational efficiency. Transformers excel overall, confirming their state-of-the-art status, although this performance comes at a higher computational cost due to the quadratic complexity of attention with respect to the sequence length. Traditional reservoirs, while inherently limited in their ability to capture complex long-range dependencies, nevertheless exhibit unexpectedly competitive results considering their simplicity and low computational overhead. Attention-enhanced reservoirs strike an intermediate balance, significantly improving over classical reservoirs while approaching transformer-level performance with adjustable complexity that depends linearly on sequence length.
We additionally extended the analysis to a large-scale token-level language modeling setting using the proposed Layered Attention-Enhanced Reservoir Computer (LAERC). By interleaving multiple fixed recurrent reservoirs with lightweight gating and feed-forward refinement, the layered architecture substantially improves over the single-reservoir setup enabling scaling to larger models. In particular, empirical scaling-law analysis confirms that the layered reservoir models follow a power-law trend in performance as a function of the number of trainable parameters. Though numerically weaker compared to the transformer, the LAERC enables the temporal mixing layer to be implemented with any physical substrate as reservoir, which would substitute the attention layer in a classical transformer.
VI Data Availability
The code used for the Shakespeare simulations is available at https://github.com/fekoester/Shakespeare_Res and the implementation of the layered reservoir language model for token-level LLM training is provided at https://github.com/fekoester/LAERC.
@@ Acknowledgements.
This study was supported in part by JSPS KAKENHI (JP22H05195, JP24KF0179, JP25H01129) and JST CREST (JPMJCR24R2) in Japan.
Appendix A Appendix: LAERC technical specification and training details
This appendix provides a self-contained technical definition of the layered reservoir language model (LAERC) used in Sec. IV and summarizes the exact training and implementation settings used for the OpenWebText simulations.
A.1 Model definition
Let $V$ denote the vocabulary size, $D$ the embedding dimension, $L$ the number of stacked LAERC blocks, and $R$ the reservoir dimension. We consider a token sequence $X=(x_{1},\dots,x_{T})$ with $x_{t}\in\{1,\dots,V\}$.
Tokens are embedded via an embedding matrix $\mathbf{E}\in\mathbb{R}^{D\times V}$,
$\mathbf{x}_{t}^{(1)}=\mathbf{E}\,\mathbf{e}(x_{t})\in\mathbb{R}^{D},$ (10)
where $\mathbf{e}(x_{t})$ is the one-hot encoding. We use weight tying, i.e., the output projection shares parameters with the input embedding: $\mathbf{W}_{\mathrm{out}}=\mathbf{E}^{\top}$.
For layer $\ell\in\{1,\dots,L\}$ and time step $t$, the block input is $\mathbf{x}_{t}^{(\ell)}\in\mathbb{R}^{D}$. Each block combines (i) a fixed recurrent reservoir path, (ii) an input-dependent mixing gate, and (iii) a feed-forward refinement, consistent with Fig. 5(b).
(i) Fixed reservoir path
The incoming representation is first normalized:
$\tilde{\mathbf{x}}_{t}^{(\ell)}=\mathrm{LN}_{\mathrm{res}}^{(\ell)}(\mathbf{x}_{t}^{(\ell)})\in\mathbb{R}^{D}.$ (11)
A single-layer tanh RNN reservoir (hidden size $R$) evolves as
$\mathbf{s}_{t}^{(\ell)}=\tanh\!\Big(\mathbf{W}_{\mathrm{res}}^{(\ell)}\mathbf{s}_{t-1}^{(\ell)}+\mathbf{W}_{\mathrm{in}}^{(\ell)}\tilde{\mathbf{x}}_{t}^{(\ell)}+\mathbf{b}^{(\ell)}\Big)\in\mathbb{R}^{R},$ (12)
where $\mathbf{W}_{\mathrm{res}}^{(\ell)}\in\mathbb{R}^{R\times R}$, $\mathbf{W}_{\mathrm{in}}^{(\ell)}\in\mathbb{R}^{R\times D}$, and $\mathbf{b}^{(\ell)}\in\mathbb{R}^{R}$.
All reservoir parameters are initialized once and then fixed (no training, and gradients are not propagated through the reservoir dynamics). The recurrent matrix is scaled so that its spectral radius satisfies
$\rho(\mathbf{W}_{\mathrm{res}}^{(\ell)})\approx\rho_{\ell}<1.$ (13)
In our simulations, we use a linear depth schedule
$\rho_{\ell}=0.95+0.04\cdot\frac{\ell-1}{\max(1,L-1)}\qquad(\ell=1,\dots,L),$ (14)
which slightly increases memory with depth.
The reservoir state is normalized prior to readout:
$\tilde{\mathbf{s}}_{t}^{(\ell)}=\mathrm{LN}_{\mathrm{res,out}}^{(\ell)}(\mathbf{s}_{t}^{(\ell)})\in\mathbb{R}^{R}.$ (15)
To map from reservoir space back to model dimension, each block contains a narrow MLP:
$\mathbf{u}_{t}^{(\ell)}$ $=\phi\!\left(\mathbf{W}^{(\ell)}_{1,\mathrm{res}}\tilde{\mathbf{s}}_{t}^{(\ell)}+\mathbf{b}^{(\ell)}_{1,\mathrm{res}}\right)\in\mathbb{R}^{H_{\mathrm{res}}},$ (16)
$\mathbf{r}_{t}^{(\ell)}$ $=\exp(\beta^{(\ell)})\left(\mathbf{W}^{(\ell)}_{2,\mathrm{res}}\mathbf{u}_{t}^{(\ell)}+\mathbf{b}^{(\ell)}_{2,\mathrm{res}}\right)\in\mathbb{R}^{D},$ (17)
where $\phi$ is GELU, $H_{\mathrm{res}}=\alpha_{\mathrm{res}}R$ is the hidden width multiplier, and $\beta^{(\ell)}$ is a learned scalar controlling the reservoir-path strength. We initialize $\beta^{(\ell)}\approx 0$ so that the reservoir readout enters at unit scale while remaining trainable.
(ii) Input-dependent mixing gate
The block combines the original representation $\mathbf{x}_{t}^{(\ell)}$ and reservoir-refined representation $\mathbf{r}_{t}^{(\ell)}$ via a feature-wise sigmoid gate:
$\mathbf{z}_{t}^{(\ell)}$ $=[\mathbf{x}_{t}^{(\ell)};\mathbf{r}_{t}^{(\ell)}]\in\mathbb{R}^{2D},$ (18)
$\hat{\mathbf{z}}_{t}^{(\ell)}$ $=\mathrm{LN}_{\mathrm{mix}}^{(\ell)}(\mathbf{z}_{t}^{(\ell)}),$ (19)
$\mathbf{g}_{t}^{(\ell)}$ $=\sigma\!\left(\mathbf{W}_{g}^{(\ell)}\hat{\mathbf{z}}_{t}^{(\ell)}+\mathbf{b}_{g}^{(\ell)}\right)\in\mathbb{R}^{D},$ (20)
$\mathbf{m}_{t}^{(\ell)}$ $=\mathbf{g}_{t}^{(\ell)}\odot\mathbf{r}_{t}^{(\ell)}+\left(\mathbf{1}-\mathbf{g}_{t}^{(\ell)}\right)\odot\mathbf{x}_{t}^{(\ell)}\in\mathbb{R}^{D}.$ (21)
(iii) Feed-forward refinement
A standard position-wise feed-forward network refines the mixed representation:
$\hat{\mathbf{m}}_{t}^{(\ell)}=\mathrm{LN}_{\mathrm{ffn}}^{(\ell)}(\mathbf{m}_{t}^{(\ell)}),$ (22)
$\mathbf{h}_{t}^{(\ell)}=\phi\!\left(\mathbf{W}^{(\ell)}_{1,\mathrm{ffn}}\hat{\mathbf{m}}_{t}^{(\ell)}+\mathbf{b}^{(\ell)}_{1,\mathrm{ffn}}\right)\in\mathbb{R}^{H_{\mathrm{ffn}}},$ (23)
$\mathbf{f}_{t}^{(\ell)}=\mathbf{W}^{(\ell)}_{2,\mathrm{ffn}}\mathbf{h}_{t}^{(\ell)}+\mathbf{b}^{(\ell)}_{2,\mathrm{ffn}}\in\mathbb{R}^{D},$ (24)
$\mathbf{x}_{t}^{(\ell+1)}=\mathbf{m}_{t}^{(\ell)}+\alpha^{(\ell)}\mathbf{f}_{t}^{(\ell)}.$ (25)
We initialize $\alpha^{(\ell)}=0$, making each block close to the identity map at initialization and improving stability for deep stacks.
After $L$ blocks, a final LayerNorm and tied projection produce logits:
$\mathbf{z}_{t}=\mathrm{LN}_{f}(\mathbf{x}_{t}^{(L+1)}),$ (26)
$\boldsymbol{\ell}_{t}=\mathbf{E}^{\top}\mathbf{z}_{t}\in\mathbb{R}^{V},$ (27)
$p(y_{t}=v\mid X)=\mathrm{softmax}(\boldsymbol{\ell}_{t})_{v}.$ (28)
Training minimizes standard next-token cross-entropy over the tokenized corpus.
A.2 Training data and batching
We train on an OpenWebText-derived corpus tokenized with a GPT-2-style BPE vocabulary (V=50,257V=50{,}257) and fixed context length $T=512$. The dataset is stored as contiguous memory-mapped integer token streams. Each microbatch is generated by sampling $B$ random start indices uniformly from the training stream, extracting contiguous chunks of length $T$, and forming next-token prediction pairs:
$\mathbf{x}=(x_{1},\dots,x_{T-1}),\qquad\mathbf{y}=(x_{2},\dots,x_{T}).$ (29)
A.3 Hyperparameters and implementation details
Unless otherwise stated, we use $R=4D$ and a linear spectral-radius schedule $\rho_{\ell}\in[0.95,0.99]$ across depth. The reservoir readout MLP width is set to $H_{\mathrm{res}}=\alpha_{\mathrm{res}}R$ with $\alpha_{\mathrm{res}}=1.0$.
We train with AdamW using $(\beta_{1},\beta_{2})=(0.9,0.999)$. We apply weight decay ($0.1$) to linear weights and use no weight decay for biases, LayerNorm parameters, and embedding parameters.
We report the effective global batch size as
$B_{\mathrm{eff}}=B\times\mathrm{accum},$ (30)
where $B=32$ is the microbatch size and $\mathrm{accum}=4$ is the number of gradient accumulation steps. For the used context length of $L=512$ thus 65536 tokens per gradient steps are used.
Training uses mixed precision on GPUs (bfloat16 when supported, otherwise float16 with dynamic loss scaling). We enable TF32 matrix multiplications on Ampere-class GPUs and cuDNN benchmarking for kernel selection. Reservoir parameters are stored in low precision when possible and remain fixed throughout training.
Embedding and linear weights are initialized from a normal distribution with standard deviation $0.02$. Output projection is tied to the embedding matrix.